Grade 12th passGeneral Physics

Find out no. Of moles of KMnO4 needed to oxidise one mole of Cu2S in acidic medium.The reaction. ;: KMnO4 + Cu2S = Mn+2 + Cu+2 + SO2. (please write the balanced reaction of Cu2S.. What should i take the vakency of Cu) ?

Parul Rajput
9 Years agoGrade 12th pass

2 Answers

Vikas TU
ApprovedApproved Tutor Answer9 Years ago
To do this U need the valemce factor not the oxidation state.
First find the valence factor for Cu in the compund using oxidation no. in both sides from the Reactant and the Product respectively.
Therfore, v.f. for Mn  => +5
and v.f. for Cu => +2 as there is no change in its valency we take the v.f as its highest state.
Therfore moles of KMnO4 x 5 = moles Cu2+ * 2
Kushagra Madhukar
5 Years ago
Dear student,
Please find the solution to your problem below.

From law of equivalence,
equivalents of Cu2S = equivalents of KMnO4
moles of Cu2S × valence factor of Cu2S = moles of KMnO4 × valence factor of KMnO4.

Now, for Cu2S we can condense the equation as
Cu+12S-2 → 2Cu2+ + S+4O4
Hence vfCu2S = 2(2 – 1) + 1(4 – ( – 2)) = 2 + 6 = 8

For KMnO4 we can write
KMn+7O4 → Mn+2 (here we haven’t considered O and K because their oxidation state remains unchanged in the reaction)
Hence vfKMnO4 = 7 – 2 = 5

Moles of Cu2S = 1
Therefore,
1 × 8 = moles of KMnO4 × 5
⇒ moles of KMnO4 = 8/5 = 1.6 moles

Hope it helps.
Thanks for the question.
Regards,
Kushagra